OpenStudy (anonymous):

Multiply and Simplify. √9x^2y * √3xy

OpenStudy (gorv):

\[\sqrt{9*x^{2}*y*3*x*y}\]

OpenStudy (gorv):

if same variable or number are multiplied than their power will be added like \[a*a=a ^{1+1}=a ^{2}\]

OpenStudy (gorv):

@moneylove01

OpenStudy (anonymous):

okay?

OpenStudy (gorv):

got it ???

OpenStudy (anonymous):

So would it be 3xy√3x

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

yeah ...great work buddy...u did it man

OpenStudy (gorv):

@moneylove01